what is the molal concentration of the solution by dissolving 2.69g of Phenol in 200 g of water
1
Expert's answer
2020-10-05T14:22:39-0400

Solution:

Cm is the molal concentration expressed in moles of solute dissolved per kilogram of solvent (mol/kg). 

Cm = (msolute / msolvent) × (1 / MW), where:

msolute is the mass of solute in grams (g),

msolvent is mass of solvent in kilograms (kg),

MW is the molecular weight of solute in g/mol. 


Solute = phenol = C6H5OH

Solvent = water = H2O


The molar mass of phenol is 94.11 g/mol.


Thus:

Cm = (2.69 g / 0.2 kg) × (1 / 94.11 g/mol) = 13.45 × 0.01063 = 0.1429 mol/kg = 0.143 m

Cm = 0.143 m


Answer: 0.143 m is the molal concentration of the solution.